Homework 6 Begin Date: 6/25/2016 12:00:00 AM -- Du
Answer #1

centrifugal force on proton Fc= mv2/r

lorenz magnetic force F = qv B

Fc= F => r = mv/(qB)

Energy of moving proton = (1/2)*mv2 = 9.2 MeV = 9.2 x10^6 x 1.6x10^-19 J= 14.72 x 10^-13 J

v2 =(2x14.72 x 10^-13)/1.67 × 10-27 = 17.63 x10^14

v=4.2x10^7 m/s

r =(1.67 × 10-27x4.2x10^7)/(1.6x10^-19 x 2.1)

radius of curvature = 0.20 m =20 cm